L^{A}TEX Quick Reference for Math Symbols Provided by the 15-251 TAs For some of these you need to do \usepackage{amssymb}.

Greek Letters

α \alpha  β \beta  γ \gamma δ \delta  ϵ \epsilon ε \varepsilon
ζ \zeta  η \eta  θ \theta ϑ \vartheta ι \iota  κ \kappa
λ \lambda μ \mu  ν \nu  ξ \xi  o o  π \pi
ϖ \varpi  ρ \rho  ϱ \varrho σ \sigma  ς \varsigma τ \tau
υ \upsilon ϕ \phi  φ \varphi χ \chi  ψ \psi  ω \omega
Γ \Gamma  Δ \Delta  Θ \Theta Λ \Lambda  Ξ \Xi  Π \Pi
Σ \Sigma  Υ \Upsilon Φ \Phi  Ψ \Psi  Ω \Omega

Binary Operations

± \pm  ∓ \mp  × \times ÷ \div  ∗ \ast
⋆ \star  ∘ \circ  • \bullet · \cdot  ∩ \cap
∪ \cup  ⊎ \uplus  ⊓ \sqcap ⊔ \sqcup  ∨ \vee
∧ \wedge  \ \setminus ≀ \wr  ⋄ \diamond      \lhd
△ \bigtriangleup        \rhd           \unlhd       \unrhd  ⊕ \oplus
▽ \bigtriangledown ⊖ \ominus  ⊗ \otimes ⊘ \oslash ⊙ \odot
◁ \triangleleft  ◯ \bigcirc † \dagger ‡ \ddagger ⨿ \amalg
▷ \triangleright

Relation Symbols

≤ \leq  ≥ \geq  ≡ \equiv  ⊧ \models  ≺ \prec  ≻ \succ
∼ \sim  ⊥ \perp  ⪯ \preceq  ⪰ \succeq  ≃ \simeq  | \mid
≪ \ll  ≫ \gg  ≍ \asymp  ∥ \parallel ⊂ \subset  ⊃ \supset
≈ \approx  ▷◁ \bowtie  ⊆ \subseteq ⊇ \supseteq  ≅ \cong            \Join
⊏ \sqsubset ⊐ \sqsupset ≠ \neq  ⌣ \smile  ⊑ \sqsubseteq ⊒ \sqsupseteq
≐ \doteq  ⌢ \frown  ∈ \in  ∋ \ni  ∝ \propto  = =
⊢ \vdash  ⊣ \dashv  < <  > >

Arrow Symbols

←  \leftarrow  ⇐ \Leftarrow  →  \rightarrow
⇒  \Rightarrow  ↔ \leftrightarrow ⇔  \Leftrightarrow
↦  \mapsto  ⟵ \longleftarrow  ⟸ \Longleftarrow
⟶ \longrightarrow  ⟹ \Longrightarrow ⟷ \longleftrightarrow
⟺ \Longleftrightarrow ⟼ \longmapsto  ↑  \uparrow
⇑  \Uparrow  ↓  \downarrow  ⇓  \Downarrow
↕  \updownarrow  ⇕  \Updownarrow  ↗  \nearrow
↘  \searrow  ↙ \swarrow  ↖  \nwarrow
↩ \hookleftarrow  ↪ \hookrightarrow ↼  \leftharpoonup
⇀  \rightharpoonup  ↽ \leftharpoondown ⇁  \rightharpoondown

Miscellaneous

. . . \ldots  · · · \cdots     ⋮ \vdots  ... \ddots  ℵ \aleph
′  \prime  ∀ \forall  ∞ \infty  ℏ \hbar  ∅ \emptyset
∃ \exists ∇ \nabla  √ \surd              \Box  △ \triangle
     \Diamond ı  \imath  ȷ \jmath  ℓ \ell  ≠ \neq
⊤ \top  ♭ \flat  ♮ \natural  ♯ \sharp  ℘ \wp
⊥ \bot  ♣ \clubsuit ♢ \diamondsuit ♡ \heartsuit ♠ \spadesuit
℧ \mho  ℜ \Re  ℑ \Im  ∠ \angle  ∂ \partial

Sum Like Symbols

∑ \sum  ∏ \prod  ∐ \coprod  ∫ \int  ∮ \oint
⋂ \bigcap ⋃ \bigcup  ⨆ \bigsqcup ⋁ \bigvee  ⋀ \bigwedge
⨀ \bigodot ⨂ \bigotimes ⨁ \bigoplus ⨄ \biguplus

Delimiters

⌊ \lfloor ⌋ \rfloor ⌈ \lceil ⌉ \rceil ⟨ \langle
⟩ \rangle | |  ∥ \|  { \{  } \}